OpenStudy (anonymous):

verify the identity (1-sinx)/cosx) = (cosx/1+sinx)

OpenStudy (luigi0210):

multiply the left-hand by 1+sinx

OpenStudy (anonymous):

So cosx=cosx?

OpenStudy (tkhunny):

Keep in mind this is valid ONLY for \(1+\sin(x) \ne 0\).

OpenStudy (luigi0210):

When you are proving an identity you only work with one side.

OpenStudy (loser66):

I am stuck, how to prove? everybody?

OpenStudy (loser66):

@Luigi0210 @tkhunny

OpenStudy (luigi0210):

Do I work it out?

OpenStudy (loser66):

not that way, we have to start from valid logic.

OpenStudy (loser66):

@dan815

OpenStudy (luigi0210):

\[\frac{ 1-sinx }{ cosx }=\frac{ cosx }{ 1+sinx }\] \[\frac{ 1-sinx }{ cosx }*\frac{ 1+sinx }{ 1+sinx }=\] \[\frac{ 1-\sin^2x }{ cosx(1+sinx) }=\] \[\frac{ \cos^2x }{ (cosx)(1+sinx) }=\] \[\frac{ cosx }{ 1+sinx }=\frac{ cosx }{ 1+sinx }\]

OpenStudy (loser66):

your logic is perfect under the condition which tkhuny indicated above. However, It cannot be used when you don't know whether it's valid or not.

OpenStudy (loser66):

this is a Prove problem, with an invalid logic, I dare not apply.